If F = (5xy --- 6x2) i + (2y - 4x) j , evaluate J F dr along the curve C in the xy plane, y = x 3 from the C point (1,1) to (2,8).

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers
Masengeli Joshua
2 years ago
r=xi+yj
dr=dxi+dyj
F.dr=(5xy-6x²)dx+(2y-4x)dy
Since y=x³
dy=3x²dx
Integrate f.dr=(5x.x³-6x²)dx+(2.x³-4x).3x²dx
|(5x⁴-6x²+6x^5-12x³)|from x=1 to x=0 
=35